### v1

# Filtrations on derived category of twisted K3 surfaces

Abstract.
In this paper, we study the Shen-Yin-Zhao type filtration on the derived category of twisted K3 surfaces. We introduce the so-called twisted Beauville-Voisin class on a twisted K3 surface and use it to extend the work in [20, 21] to the case of twisted K3 surfaces. It shares many nice properties as the untwisted case. For instance, the filtration is preserved under any Fourier-Mukai transformation. As an application, we obtain a new filtration on the Chow group of zero cycles on the Bridgeland moduli space of twisted K3 surfaces. It is birational invariant and can be viewed as a candidate for the conjectural Beauville-Voisin filtration. We compare this filtration with Voisin’s filtration and show they coincide on the locally free loci. Some other applications (such as Bloch’s conjecture for birational automorphisms) are discussed at the end.
Key words and phrases:
twisted K3 surfaces, derived category, Hyper-Kähler varieties, Beauville-Voisin filtration, Bloch’s conjecture
2020 Mathematics Subject Classification:
14C25, 14F08, 14J28, 14J42
1. Introduction
1.1. Filtrations on K3 category
Let $X$ be a K3 surface and let $\mathfrak{o}_{X}\in\operatorname{CH}_{0}(X)$ be the Beauville-Voisin class defined in [3]. In [20], O’Grady introduced a filtration $\mathbf{S}_{\bullet}(X)$ on $\operatorname{CH}_{0}(X)$ given by
$\mathbf{S}_{i}(X)=\bigcup_{\deg([z])=i}\Big{\{}[z]+\mathbb{Z}\cdot[\mathfrak{o}_{X}]\Big{\}}$
for all effective zero-cycles $z$ of degree $i$. Denote by $\mathrm{D}^{b}(X)$ the bounded derived category of coherent sheaves on $X$. In [21], Shen-Yin-Zhao used O’Grady’s filtration to construct a natural increasing filtration on the bounded derived category $\mathrm{D}^{b}(X)$
$\mathbf{S}_{0}(\mathrm{D}^{b}(X))\subset\mathbf{S}_{1}(\mathrm{D}^{b}(X))\subset\ldots\subset\mathbf{S}_{i}(\mathrm{D}^{b}(X))\ldots$
where $\mathbf{S}_{i}(\mathrm{D}^{b}(X))$ consists of objects $\mathcal{E}\in\mathrm{D}^{b}(X)$ with $c_{2}(\mathcal{E})\in\mathbf{S}_{i}(X)$. They have shown that for an object $\mathcal{E}\in\mathrm{D}^{b}(X)$, one has
$c_{2}(\mathcal{E})\in\mathbf{S}_{\mathrm{d}(\mathcal{E})}(X),$
where $d(\mathcal{E})=\frac{1}{2}\dim\operatorname{Ext}^{1}(\mathcal{E},\mathcal{E})$. More generally, let $\mathscr{X}\to X$ be a twisted K3 surface in the sense of [16]. Denote by $\mathrm{D}^{b}(\mathscr{X}):=\mathrm{D}^{(1)}(\mathscr{X})$ the bounded derived category of 1-fold twisted coherent sheaves on $\mathscr{X}\to X$ (See §2.1). It is expected that there is a similar filtration on $\mathrm{D}^{b}(\mathscr{X})$. The main result of this paper is
Theorem 1.2.
There exists a unique cycle class $\mathfrak{o}_{\mathscr{X}}\in\operatorname{CH}_{0}(X)$ of degree $1$ such that for any $\mathcal{E}\in\mathrm{D}^{b}(\mathscr{X})$, we have
$c_{2}(\mathcal{E})-\operatorname{rank}(\mathcal{E})\mathfrak{o}_{\mathscr{X}}\in\mathbf{S}_{\mathrm{d}(\mathcal{E})}(X),$
where $\mathrm{d}(\mathcal{E})=\frac{1}{2}\operatorname{Ext}^{1}(\mathcal{E},\mathcal{E})$.
We may call $\mathfrak{o}_{\mathscr{X}}$ the $\mathscr{X}$-twisted Beauville-Voisin class on $X$. From our construction, when $\mathscr{X}\to X$ is essentially trivial, it can be easily shown that $\mathfrak{o}_{\mathscr{X}}=\mathfrak{o}_{X}$. One may expect that $\mathfrak{o}_{\mathscr{X}}=\mathfrak{o}_{X}$ for any $\mathscr{X}\to X$. An evidence is that if $X$ is defined over $\bar{\mathbb{Q}}$, one can show that $\mathfrak{o}_{\mathscr{X}}$ is also defined over $\bar{\mathbb{Q}}$ (See Remark 4.8) and thus $\mathfrak{o}_{\mathscr{X}}=\mathfrak{o}_{X}$ by the Bloch-Beilinson conjecture.
There are some useful consequences of Theorem 1.2. For instance, one can obtain a generalization of [9, Corollary 2.6]
Corollary 1.3.
For two rigid objects $\mathcal{E},\mathcal{F}\in\mathrm{D}^{b}(\mathscr{X})$, $c_{2}(\mathcal{E})=c_{2}(\mathcal{F})\in\operatorname{CH}_{0}(X)$ if they have the same Mukai vector.
We shall mention that for in general $\mathrm{D}^{b}(\mathscr{X})$ may not have rigid objects. Moreover, motivated from the work of [21], one can define a filtration on $\mathrm{D}^{b}(\mathscr{X})$ as follows:
$\mathbf{S}_{i}(\mathrm{D}^{b}(\mathscr{X}))=\Big{\{}\mathcal{E}\in\mathrm{D}^{b}(\mathscr{X})|~{}c_{2}(\mathcal{E})-\operatorname{rank}(\mathcal{E})\mathfrak{o}_{\mathscr{X}}\in\mathbf{S}_{i}(X)\Big{\}}.$ (1)
Then we have
Theorem 1.4.
For any derived equivalence $\Psi:\mathrm{D}^{b}(\mathscr{X})\to\mathrm{D}^{b}(\mathscr{Y})$ between two twisted K3 surfaces $\mathscr{X}\to X$ and $\mathscr{Y}\to Y$, we have
$\mathcal{E}\in\mathbf{S}_{i}(\mathrm{D}^{b}(\mathscr{Y}))\Leftrightarrow\Psi(\mathcal{E})\in\mathbf{S}_{i}(\mathrm{D}^{b}(\mathscr{X})).$
In other words, the filtration (1) is preserved under derived equivalences.
See [21, Corollary 2] for the untwisted case.
1.5. BV filtration on Bridgeland moduli spaces
The proof of Theorem 1.2 and Theorem 1.4 involves a dedicated study of so-called Beauville-Voisin filtration on moduli space of stable objects in $\mathrm{D}^{b}(\mathscr{X})$. Let $\mathbf{v}\in\widetilde{\mathrm{H}}(\mathscr{X})$ be a primitive Mukai vector and $\sigma$ be a $\mathbf{v}$-generic stability condition on $\mathrm{D}^{b}(\mathscr{X})$. The (coarse) moduli space $M=M_{\sigma}(\mathscr{X},\mathbf{v})$ of $\sigma$-stable objects in $\mathrm{D}^{b}(\mathscr{X})$ with Mukai vector $\mathbf{v}$ is a smooth projective hyper-Kähler variety. Voisin has introduced a filtration $\mathbf{S}_{\bullet}^{\mathrm{B}\mathrm{V}}\operatorname{CH}_{0}(M)\subseteq\operatorname{CH}_{0}(M)$ given by
$\mathbf{S}_{i}^{\rm BV}\operatorname{CH}_{0}(M):=\left<x\in M|~{}\dim O_{x}\geq d-i\right>$
where $O_{x}=\{x^{\prime}\in M|~{}[x]=[x^{\prime}]\in\operatorname{CH}_{0}(M)\}$. From the definition, one can see such filtration only depends on the isomorphic class of $M$. On the other hand, there is a natural filtration
$\mathbf{S}^{\rm SYZ}_{\bullet}\operatorname{CH}_{0}(M)\subseteq\operatorname{CH}_{0}(M)$
given by the restriction of the filtration $\mathbf{S}_{\bullet}(\mathrm{D}^{b}(\mathscr{X}))$ to $\operatorname{CH}_{0}(M)$, i.e.
$\mathbf{S}^{\rm SYZ}_{i}\operatorname{CH}_{0}(M)=\left<x\in M|~{}x\in\mathbf{S}_{i}(\mathrm{D}^{b}(\mathscr{X}))\right>$ (2)
Then Theorem 1.4 actually says that filtration $\mathbf{S}^{\rm SYZ}_{\bullet}\operatorname{CH}_{0}(M)$ does not depend on the moduli interpretation of $M$. A natural expectation is that
$\mathbf{S}^{\rm SYZ}_{\bullet}\operatorname{CH}_{0}(M)=\mathbf{S}^{\rm BV}_{\bullet}\operatorname{CH}_{0}(M),$
just as the untwisted case (cf[15, Theorem 1.1]). Another goal of this paper is to verify this for arbitrary $M$. The main technical result is
Theorem 1.6.
Assume $\mathbf{v}^{2}=2n-2\geq 0$. Let $M^{\rm l.f.}\subseteq M$ be the open locus (if non-empty) consisting of $\sigma$-stable locally free 1-fold $\mathscr{X}$-twisted sheaves. Then
$\dim\mathcal{O}_{\mathcal{E}}\cap M^{\rm l.f.}\geq n-i\Leftrightarrow\mathcal{E}\in\mathbf{S}_{i}(\mathrm{D}^{b}(\mathscr{X})).$
In particular, $\mathbf{S}_{i}^{\operatorname{BV}}\operatorname{CH}_{0}(M)=\mathbf{S}_{i}^{\rm SYZ}\operatorname{CH}_{0}(M)$ if $M=M^{\rm l.f.}$.
This is proved by studying the degenerate loci of morphisms between twisted sheaves, which generalizes the work in [26] and [15]. Our method can be also applied to the case of hyper-Kähler varieties of generalized Kummer type (see some discussions in §4.15). According to the work in [7, Theorem 3.4], the locally freeness assumption can be released once $\mathscr{X}$ is derived equivalent to an untwisted K3 surface (cf[15, Proof of Theorem 1.1]). Roughly speaking, it has been shown that in that case $M$ is isomorphic to a Bridgeland moduli space which consists of slope stable locally free sheaves. This yields
Corollary 1.7.
If $\mathrm{D}^{b}(\mathscr{X})\cong\mathrm{D}^{b}(Y)$ for some K3 surface $Y$, then $\mathbf{S}_{i}^{\rm BV}\operatorname{CH}_{0}(M)=\mathbf{S}_{i}^{\rm SYZ}\operatorname{CH}_{0}(M)$
The existence of such filtration plays an important role in the study of Bloch’s conjecture for zero cycles on hyper-Kähler varieties. An advantage of Shen-Yin-Zhao type filtration $\mathbf{S}_{\bullet}^{\rm SYZ}\operatorname{CH}_{0}$ is that it is preserved under birational transformations and the graded pieces are controlled by the first one.
Theorem 1.8.
Let $Y$ be a smooth projective hyper-Kähler variety of $K3^{[n]}$-type with a birational Lagrangian fibration. Then $\operatorname{CH}_{0}(Y)$ can be equipped with a natural filtration in (2) and Bloch’s conjecture holds for any (anti-)symplectic birational automorphism preserving the birational Lagrangian fibration, i.e.
$\phi^{\ast}|_{\mathrm{H}^{2,0}(Y)}=\pm\operatorname{id}\Leftrightarrow\phi_{\ast}|_{\mathrm{Gr}_{s}\operatorname{CH}_{0}(Y)}=(\pm 1)^{s}\operatorname{id},~{}\forall~{}1\leq s\leq n.$ (3)
where $\mathrm{Gr}_{s}\operatorname{CH}_{0}(Y)$ is the $s$-th graded piece of the the filtration $\mathbf{S}_{\bullet}^{\rm SYZ}\operatorname{CH}_{0}(Y)$.
Another application is concerned with the existence of algebraically coisotropic subvarieties. With the same argument in [21, Theorem 0.5], one can obtain
Corollary 1.9.
For $0\leq i\leq\frac{\mathbf{v}^{2}+2}{2}$, there exist algebraically coisotropic subvarieties
$\iota:Z_{i}\dashrightarrow B_{i}$
where $Z_{i}\subseteq M$ has codimension $i$ and the general fibers of $\iota$ are constant cycle subvarieties of $M$ with dimension $i$.
1.10. Organization of the paper
In Section 2, we review the basic results on twisted K3 surfaces and their derived categories. We discuss the twisted Chern character and its connection to the Chern character on gerbes. In Section 3, we study filtrations on the moduli space $M_{H}(\mathscr{X},\mathbf{v})$ of Gieseker stable $\mathscr{X}$-twisted torsion sheaves. Such a moduli space admits a natural Lagrangian fibration. The key observation is that there exists a finite incidence correspondence between the smooth support loci of $M_{H}(\mathscr{X},\mathbf{v})$ for various $\mathscr{X}\to X$. Section 4 is the main part of this paper. We construct the twisted Beauville-Voisin class via using the Chern class of stable twisted locally free sheaves with maximal dimensional rational equivalent orbit. By investigating the degenerate loci of morphisms between polystable twisted locally free sheaves, we show our twisted BV class is well-defined and prove Theorem 1.2. Section 5 and Section 6 are devoted to the proof of Theorem 1.4 and Theorem 1.8.
Acknowledgement
We want to thank Ziyu Zhang and Junliang Shen for helpful discussions. The first author is supported by NSFC grant (No. 12121001, No. 12171090) and Shanghai Pilot Program for Basic Research (No. 21TQ00).
Convention
In this paper, we use $\operatorname{CH}^{\bullet}(X)$ to denote the Chow groups of $X$ with rational coefficients.
2. Preliminary
Throughout this section, we work over $\mathbb{C}$.
2.1. Twisted K3 surface and twisted sheaves
Let $p:\mathscr{X}\rightarrow X$ be a $\mu_{m}$-gerbe over a smooth projective K3 surface $X$. This corresponds to a pair $(X,\alpha)$ for some $\alpha\in\mathrm{H}^{2}_{\operatorname{fl}}(X,\mu_{m})$, where the cohomology group is with respect to the flat topology. Recall that there is a Kummer exact sequence
$1\rightarrow\mu_{m}\rightarrow{\mathcal{O}_{X}^{\times}}\xrightarrow{x\mapsto x^{m}}{\mathcal{O}_{X}^{\times}}\rightarrow 1$ (4)
in flat topology and it induces a surjective map
$\mathrm{H}^{2}_{\rm fl}(X,\mu_{m})\rightarrow\mathrm{Br}(X)[m].$ (5)
We denote by $[\mathscr{X}]$ the image of $\alpha$ in $\operatorname{Br}(X)[m]$. We say $\mathscr{X}\to X$ is essentially trivial if the associated Brauer class $[\mathscr{X}]$ is zero in $\operatorname{Br}(X)$.
A $d$-fold $\mathscr{X}$-twisted sheaf on $p:\mathscr{X}\to X$ is an $\mathcal{O}_{\mathscr{X}}$-module of weight $d$ with respect to the $\mu_{m}$-action (cf[16, Def 2.1.2.2]). Let $\mathrm{D}^{(d)}(\mathscr{X})$ be the bounded derived category of coherent $d$-fold $\mathscr{X}$-twisted sheaves. For a $\mathscr{X}$-twisted locally free sheaf $\mathcal{E}\in\mathrm{D}^{(1)}(\mathscr{X})$, we have a diagram
[…]
where $\mathbb{P}(\mathcal{E})\to Y$ is essentially trivial and $Y$ is served as a Severi-Brauer variety of $(X,\alpha)$.
Moreover, there is a derived tensor product
$-\otimes^{\mathbb{L}}-:\mathrm{D}^{(d)}(\mathscr{X})\times\mathrm{D}^{(d^{\prime})}(\mathscr{X})\to\mathrm{D}^{(d+d^{\prime})}(\mathscr{X})$
and the derived dual functor
$(-)^{\vee}:\mathrm{D}^{(d)}(\mathscr{X})\to\mathrm{D}^{(-d)}(\mathscr{X})^{\rm opp}.$
When $d=0$, one may identify $\mathrm{D}^{(0)}(\mathscr{X})=\mathrm{D}^{b}(X)$ via the pullback functor.
2.2. Chern character of twisted sheaves
Let $K_{0}(\mathscr{X})$ be the Grothendieck group of $1$-fold $\mathscr{X}$-twisted coherent sheaves. In [17, §3.3.4], Lieblich-Maulik-Snowden have introduced a twisted Chern character
$\operatorname{ch}_{\mathscr{X}}:K_{0}(\mathscr{X})\to\operatorname{CH}^{\ast}(X),$ (6)
which takes values in the Chow ring of $X$. For a locally free sheaf $\mathcal{F}$, it is defined by
$\operatorname{ch}_{\mathscr{X}}(\mathcal{F})=\sqrt[m]{\operatorname{ch}(p_{\ast}\mathcal{F}^{\otimes m})}$ (7)
where the $m$-th root is uniquely determined by requiring $r(\operatorname{ch}_{\mathscr{X}}(\mathcal{F}))=\operatorname{rank}(\mathcal{F})$. One can define the twisted Chern class
$c_{i}:K_{0}(\mathscr{X})\to\operatorname{CH}^{i}(X)$
accordingly. The definition above avoids the use of the Chow rings of stacks. However, one can compare it with the Chern character $\operatorname{ch}:K_{0}(\mathscr{X})\to\operatorname{CH}^{\ast}(\mathscr{X})$ for stacks. An easy fact is
Lemma 2.3.
For any $x\in K_{0}(\mathscr{X})$, we have $p_{\ast}\operatorname{ch}(x)=\frac{1}{m}\operatorname{ch}_{\mathscr{X}}(x)$.
Proof.
It suffices to check the case when $x$ is the class of a locally free sheaf $\mathcal{F}$.
According to [23], the pullback map
$p^{\ast}:\operatorname{CH}^{\ast}(X)\to\operatorname{CH}^{\ast}(\mathscr{X})$
is a ring isomorphism. Hence there is a unique $z\in\operatorname{CH}^{\ast}(X)$ such that $\operatorname{ch}(\mathcal{F})=p^{*}z$. Then we have
$p_{*}(\operatorname{ch}(\mathcal{F})^{m})=p_{*}p^{*}z^{m}=\dfrac{1}{m}z^{m}.$
On the other hand, from the definition (7), we have
$p^{*}(\operatorname{ch}_{\mathscr{X}}(\mathcal{F})^{m})=p^{*}\operatorname{ch}(p_{\ast}\mathcal{F}^{\otimes m})=\operatorname{ch}(p^{*}p_{\ast}\mathcal{F}^{\otimes m})=\operatorname{ch}(\mathcal{F}^{\otimes m})=\operatorname{ch}(\mathcal{F})^{m}.$
Here we use the fact that $p^{\ast}\circ p_{\ast}$ is identity for 00-twisted sheaves. The assertion then follows from the projection formula on Chow groups. ∎
Example 2.4.
(Chern class of torsion sheaves) Let us consider the Chern class of a twisted torsion sheaf whose schematic support is an irreducible curve $C$ on $X$. There is a Cartesian diagram
[…]
Let $\mathcal{L}$ be a 1-fold $\mathscr{C}$-twisted coherent sheaf. For the $\mathscr{X}$-twisted torsion sheaf $\Phi_{\ast}\mathcal{L}$, we have
$\frac{1}{m}\operatorname{ch}_{\mathscr{X}}(\Phi_{\ast}\mathcal{L})$ $=p_{\ast}\operatorname{ch}(\Phi_{\ast}\mathcal{L})$ (8)
$=p_{\ast}(\Phi_{\ast}(\operatorname{ch}(\mathcal{L})\cdot\mathrm{td}_{\mathscr{C}/\mathscr{X}}))$
$=\phi_{\ast}q_{\ast}(\operatorname{ch}(\mathcal{L})\cdot\mathrm{td}_{\mathscr{C}/\mathscr{X}})$
$=\phi_{\ast}(q_{\ast}(\operatorname{ch}(\mathcal{L}))\cdot\mathrm{td}_{C/X})=\frac{1}{m}\phi_{\ast}\operatorname{ch}_{\mathscr{C}}(\mathcal{L})\cdot(1-\frac{1}{2}[C]).$
In particular, we have $c_{2}(\Phi_{\ast}\mathcal{L})=[C]^{2}-\phi_{\ast}(c_{1}(\mathcal{L}))\in\operatorname{CH}_{0}(X)$. It has the same form as in the untwisted case. The foregoing argument uses the Grothendieck-Riemann-Roch theorem on $\mu_{m}$-gerbes (cf[16, Proposition 2.2.7.8]). Alternatively one can obtain the above formula using the same method as [4, Lemma 4.1.4].
Similarly, we can define the topological Chern characters $\operatorname{ch}_{\mathscr{X}}^{\rm top}$ of elements in $K_{0}(\mathscr{X})$. Moreover, let us briefly review the construction of twisted Mukai lattice of $\mathscr{X}\to X$. Recall that there is a B-field
$\mathbf{B}\in\mathrm{H}^{2}_{\rm tr}(X,\mathbb{Q}):=\operatorname{NS}(X)^{\perp}$
such that $[\mathscr{X}]=\delta(B)$, where $\delta:\mathrm{H}^{2}(X,\mathbb{Q})\to\mathrm{H}^{2}(X,\mathcal{O}_{X}^{\times})$ is induced from the exponential exact sequence. We define
$\widetilde{\mathrm{H}}(X)=\mathbb{Z}\oplus\mathrm{H}^{2}(X,\mathbb{Z})\oplus\mathbb{Z}$
to be the (untwisted) Mukai lattice under the Mukai pairing $\langle$,$\rangle$ , ⟩. We use $\widetilde{\mathrm{H}}_{\rm alg}(X)=\mathbb{Z}\oplus\operatorname{NS}(X)\oplus\mathbb{Z}$ to denote the algebraic part of $\widetilde{\mathrm{H}}(X)$. There is a rational isometry given by
$\exp(\mathbf{B}):\widetilde{\mathrm{H}}(X)\otimes\mathbb{Q}$ $\longrightarrow\widetilde{\mathrm{H}}(X)\otimes\mathbb{Q}$
$(r,L,s)$ $\mapsto(r,L+rB,s+\frac{rB^{2}}{2}).$
Definition 2.1.
We define the twisted Mukai lattice of $\mathscr{X}\to X$ as
$\widetilde{\mathrm{H}}(\mathscr{X})\coloneqq\exp(\mathbf{B})\cdot\widetilde{\mathrm{H}}(X)\subset\widetilde{\mathrm{H}}(X)\otimes_{\mathbb{Z}}\mathbb{Q},$
Up to a lattice isometry, $\widetilde{\mathrm{H}}(\mathscr{X})$ is independent of the choice of a $B$-field lift. One can similarly define the twisted algebraic Mukai lattice $\widetilde{\mathrm{H}}_{\rm alg}(\mathscr{X})$ by replacing $\widetilde{\mathrm{H}}(X)$ with $\widetilde{\mathrm{H}}_{\rm alg}(X)$. For $\mathcal{F}\in\mathrm{D}^{b}(\mathscr{X})$, its Mukai vector is defined by
$\mathbf{v}^{B}(\mathcal{F})=\exp(\mathbf{B})\operatorname{ch}^{\rm top}_{\mathscr{X}}(\mathcal{F})\sqrt{\operatorname{td}_{X}^{\rm top}}\in\widetilde{\mathrm{H}}(\mathscr{X}).$
In this paper, we will fix a $B$-field $\mathbf{B}$ and simply write it as $\mathbf{v}(\mathcal{F})$.
2.5. Bridgeland moduli space
Let $\mathbf{v}=(r,L,s)\in\widetilde{\mathrm{H}}(\mathscr{X})$ be a primitive Mukai vector with $\mathbf{v}^{2}\geq 0$ and let $\sigma$ be a $\mathbf{v}$-generic Bridgeland stability condition. Let $\mathscr{M}_{\sigma}(\mathscr{X},\mathbf{v})$ be the moduli stack of $\sigma$-stable objects in $\mathrm{D}^{b}(\mathscr{X})$ with Mukai vector $\mathbf{v}$. Let $\mathscr{M}_{\sigma}(\mathscr{X},\mathbf{v})\to M_{\sigma}(\mathscr{X},\mathbf{v})$ be the coarse moduli map, which can be also viewed as a gerbe on $M_{\sigma}(\mathscr{X},\mathbf{v})$. In this way, we have a universal object
$\mathscr{E}\in\mathrm{D}^{(-1,1)}(\mathscr{X}\times\mathscr{M}_{\sigma}(\mathscr{X},\mathbf{v})),$
on $\mathscr{X}\times\mathscr{M}_{\sigma}(\mathscr{X},\mathbf{v})$.
According to [27, Sections 7 and 8], [5, Proposition 14.2] and [2, Theorem 6.10, Section 7], when $\mathbf{v}^{2}\geq-2$, $M_{\sigma}(\mathscr{X},\mathbf{v})$ is a non-empty smooth hyper-Kähler variety of K3${}^{[n]}$-type of dimension $2+v^{2}$. If $\sigma$ is a Gieseker stability condition with respect to an ample line bundle $H\in\operatorname{Pic}(X)$, we may write it as $\mathscr{M}_{H}(\mathscr{X},\mathbf{v})$. When $r\geq 0$ and $\mathbf{v}^{2}\geq-2$, the moduli space $\mathscr{M}_{H}(\mathscr{X},\mathbf{v})$ is not empty for a $\mathbf{v}$-generic $H$. There is a natural isometry
$\mathbf{v}^{\perp}\xrightarrow{\cong}\operatorname{NS}(M_{\sigma}(\mathscr{X},\mathbf{v}))$ (9)
induced by the (quasi-)universal object, where $\mathbf{v}^{\perp}$ is the orthogonal complement of $\mathbf{v}$ in $\widetilde{\mathrm{H}}_{\rm alg}(\mathscr{X})$ (cf[27, Sections 7 and 8]).
In the study of zero cycles on $M_{\sigma}(\mathscr{X},\mathbf{v})$, a very important result proved by Marian-Zhao is
Theorem 2.6 ([19]).
Let $\mathcal{E},\mathcal{F}\in M_{\sigma}(\mathscr{X},\mathbf{v})$. Then $c_{2}(\mathcal{E})=c_{2}(\mathcal{F})\in\operatorname{CH}_{0}(X)$ if and only if $[\mathcal{E}]=[\mathcal{F}]\in\operatorname{CH}_{0}(M_{\sigma}(\mathscr{X},\mathbf{v}))$.
We also need the square zero conjecture for hyper-Kähler varieties of $K3^{[n]}$-type.
Theorem 2.7.
[2, Theorem 1.1, Theorem 1.5] With notations as above. Assume that there exists an isotropic class $\mathbf{w}\in\mathbf{v}^{\perp}$, then $M_{\sigma}(\mathscr{X},\mathbf{v})$ admits a birational Lagrangian fibration.
3. Cycles on HK with birational Lagrangian fibration
In this section, we investigate the zero cycles on moduli space of twisted torsion sheaves.
3.1. Density of constant cycle Lagrangian subvarieties
Let us recall some general results for cycles on hyper-Kähler varieties with a Lagrangian fibration. The first result is a strengthen of [18, Theorem 1.3]
Proposition 3.2.
If a hyper-Kähler variety $M$ admits a birational Lagrangian fibration, then there is a dense collection of constant cycle Lagrangian subvarieties. Moreover, points on those constant cycle Lagrangian subvarieties represent the same class.
Proof.
As the property is preserved under birational transformations, we are reduced to consider the case $M$ admits a Lagrangian fibration $M\to\mathbb{P}^{n}$. If $T\to\mathbb{P}^{n}$ is a morphism, we set
$M_{T}:=M\times_{\mathbb{P}^{n}}T.$
Let $U\subseteq\mathbb{P}^{n}$ be the open locus parametrizing all the smooth fibers of $M\to\mathbb{P}^{n}$. According to [18, Theorem 1.3], there is a constant cycle Lagrangian subvariety $\Sigma\subseteq M$ such that the restriction $\Sigma_{U}=\Sigma\cap M_{U}\to U$ is finite. This gives a map
$\operatorname{Spec}(\mathbb{C}(\Sigma))\to\operatorname{Spec}(\mathbb{C}(\mathbb{P}^{n}))\to\mathbb{P}^{n}.$ (10)
Let $\eta$ be the generic point of $\mathbb{P}^{n}$ and let $M_{\eta}$ be the generic fiber of $M\to\mathbb{P}^{n}$. The multisection $\Sigma_{U}$ defines a closed point $0_{\Sigma}\in M_{\eta}(\mathbb{C}(\Sigma))$ via (10). Now we may consider the torsion points on the abelian variety $(M_{\bar{\eta}},0_{\Sigma})$. For each geometrically closed torsion point $p\in M_{\bar{\eta}}$, it can be spread out to be a multisection on the abelian fibration $M_{U}\to U$. Its Zariski closure $\mathcal{Z}_{p}\subseteq M$ is a constant cycle Lagrangian subvariety on $M$. As the geometrically closed torsion points of $M_{\bar{\eta}}$ are dense and $\operatorname{CH}_{0}(M)$ is torsion-free, the assertion follows. ∎
Remark 3.3.
From the construction, one can see that all these constant cycle Lagrangian subvarieties are contained in a single orbit $O_{x}$ with $x\in\Sigma$.
3.4. Twisted torsion sheaves with maximal dimensional rational equivalent orbit
Let $\mathbf{v}=(0,L,s)\in\widetilde{\mathrm{H}}(X)$ be a primitive vector with $L\in\operatorname{Pic}(X)$ effective. Let $H$ be $v$-generic. The moduli space $M_{H}(\mathscr{X},\mathbf{v})$ parameterizes torsion sheaves of pure dimension $1$. Their schematic support are curves in $|L|$. Moreover, if the support curve $\mathscr{C}\to C$ of $\mathcal{E}\in M_{H}(\mathscr{X},\mathbf{v})$ is integral, then $\operatorname{rank}(\mathcal{E}|_{\mathscr{C}})=1$ since $c_{1}(\mathcal{E})=[C]$.
There is a Lagrangian fibration
$\pi_{\mathscr{X}}:M_{H}(\mathscr{X},\mathbf{v})\to|L|.$ (11)
by sending $\mathcal{E}$ to its schematic support.
Proposition 3.5.
For $\mathcal{E}\in M_{H}(\mathscr{X},\mathbf{v})$, we have $c_{2}(\mathcal{E})\in\mathbf{S}_{\mathrm{d}(\mathcal{E})}(X)$.
Proof.
By Example 2.4, $c_{2}(\mathcal{E})$ is a zero cycle supported on a curve $C\in|L|$. Since the genus of $C$ is less than or equal to $\frac{1}{2}(L^{2}+2)=\mathrm{d}(\mathcal{E})$, the assertion then follows from [20, Claim 0.2] (see also [26, Proposition 15]). ∎
As in [26] and [21], set $\dim M_{\sigma}(\mathscr{X},\mathbf{v})=2n$ and we can define the incidence
$\Gamma^{\prime}:=\Big{\{}(\xi,\mathcal{E})~{}|~{}c_{2}(\mathcal{E})=[\xi]+k\mathfrak{o}_{X}\in\operatorname{CH}_{0}(X)\Big{\}}\subset X^{[n]}\times M_{\sigma}(\mathscr{X},\mathbf{v}),$ (12)
where $k\in\mathbb{Q}$ is a constant determined by $\mathbf{v}$. It is known that $\Gamma^{\prime}$ is a countable union of closed subsets of $X^{[n]}\times M_{\sigma}(\mathscr{X},\mathbf{v})$ and there are two projections
$\pi_{1}:\Gamma^{\prime}\to M_{\sigma}(\mathscr{X},\mathbf{v}),\quad\pi_{2}:\Gamma^{\prime}\to X^{[n]}.$
Due to Proposition 3.5, $\pi_{1}$ is dominant. By the same argument in [20, Proposition 1.3], we know that $\pi_{2}$ is also dominant.
By cutting $\Gamma^{\prime}$ via hyperplane sections, we can obtain a subvariety $\Gamma\subseteq\Gamma^{\prime}$ such that $\Gamma$ is generically finite over $M_{\sigma}(\mathscr{X},\mathbf{v})$ and $X^{[n]}$. Then we have
$c_{2}(\mathcal{E})\in\mathbf{S}_{i}(X)\Rightarrow\dim\mathcal{O}_{\mathcal{E}}\geq\mathrm{d}(\mathcal{E})-i.$ (13)
for any $\mathcal{E}\in M_{\sigma}(\mathscr{X},\mathbf{v})$. This yields
Corollary 3.6.
Let $\Sigma\subseteq M_{\sigma}(\mathscr{X},\mathbf{v})$ be a constant cycle Lagrangian subvariety in Proposition 3.2. Then $c_{2}(\mathcal{E})\in\mathbf{S}_{0}(X)$ for any $\mathcal{E}\in\Sigma$.
Proof.
According to the construction in the proof of Proposition 3.2, we get a Lagrangian constant cycle subvariety $\Sigma_{\mathscr{X}}^{\prime}$ such that for any $\mathcal{E}^{\prime}\in\Sigma_{\mathscr{X}}^{\prime}$, we have $c_{2}(\mathcal{E}^{\prime})=c_{2}(\mathcal{E})$.
Moreover, we have the density of such Lagrangian constant cycle subvarieties. Hence we may assume that $\dim\pi_{2}(\pi_{1}^{-1}(\Sigma_{\mathscr{X}}^{\prime}))=d(\mathcal{E})$. According to Theorem 2.6, $\pi_{2}(\pi_{1}^{-1}(\Sigma_{\mathscr{X}}^{\prime}))$ is a Lagrangian constant cycle subvariety in $X^{[n]}$.
Then for any $\mathcal{F}\in\pi_{2}(\pi_{1}^{-1}(\Sigma_{\mathscr{X}}^{\prime}))$, we have $c_{2}(\mathcal{F})\in\mathbf{S}_{0}(X)$ by [26, Theorem 9]. Therefore $c_{2}(\mathcal{E}^{\prime})\in\mathbf{S}_{0}(X)$ by the definition of $\Gamma$ for any $\mathcal{E}^{\prime}\in\Sigma_{\mathscr{X}}^{\prime}$. ∎
3.7. Correspondence between torsion sheaves
With notations as above, let us write
$M_{\mathscr{X}}:=M_{H}(\mathscr{X},\mathbf{v})$
for short. Assume $L\in\operatorname{Pic}(X)$ is effective. We denote by $\mathcal{C}\to|L|$ the complete family of curves in $|L|$ and $\mathcal{C}^{\operatorname{sm}}\to|L|^{\operatorname{sm}}$ the family of smooth curves in $|L|$. Let $M^{\operatorname{sm}}_{\mathscr{X}}=M_{\mathscr{X}}\times_{|L|}|L|^{\operatorname{sm}}$. Then $M^{\operatorname{sm}}_{\mathscr{X}}\to|L|^{\operatorname{sm}}$ is a torsor of the relative Jacobian $\operatorname{Pic}^{0}(\mathcal{C}^{\operatorname{sm}}/|L|^{\operatorname{sm}})$ with the action $M^{\operatorname{sm}}_{\mathscr{X}}\times_{|L|^{\operatorname{sm}}}\operatorname{Pic}^{0}(\mathcal{C}^{\operatorname{sm}}/|L|^{\operatorname{sm}})\to M^{\operatorname{sm}}_{\mathscr{X}}$ given by
$(\mathcal{E},L)\mapsto\mathcal{E}\otimes(i_{C})_{\ast}L$
where $\mathcal{E}\in M_{\mathscr{X}}^{\operatorname{sm}}$ with $\mathrm{supp}(\mathcal{E})=C$ and $L\in\operatorname{Pic}^{0}(C)$ (cf[11, Subsection 3.2]).
Theorem 3.8.
For $\mathcal{E}\in M^{\operatorname{sm}}_{\mathscr{X}}$, we have
$c_{2}(\mathcal{E})\in\mathbf{S}_{i}(X)\Leftrightarrow\dim(O_{\mathcal{E}}\cap M^{\operatorname{sm}}_{\mathscr{X}}(\mathbf{v}))\geq\mathrm{d}(\mathcal{E})-i.$ (14)
Proof.
If $\mathscr{X}\to X$ is essentially trivial, this is established in [15, Theorem 3.1 & Theorem 1.1]. If $\mathscr{X}\to X$ is not essentially trivial, we can establish a finite correspondence between $M_{X}^{\operatorname{sm}}$ and $M_{\mathscr{X}}^{\operatorname{sm}}$ which preserves the two filtrations. More precisely, consider the incidence variety
$R=\Big{\{}(\mathcal{E},E)\in M^{\operatorname{sm}}_{\mathscr{X}}\times M_{X}^{\operatorname{sm}}|~{}c_{2}(\mathcal{E})-c_{2}(E)\in\mathbf{S}_{0}(X)~{}\Big{\}},$ (15)
which is a countable union of closed subvarieties of $M^{\operatorname{sm}}_{\mathscr{X}}\times M_{X}^{\operatorname{sm}}$. Then
Claim 1.
There exists a subvariety $W\subseteq R$ such that the projections $\pi_{1}:W\to M_{\mathscr{X}}^{\operatorname{sm}}$ and $\pi_{2}:W\to M_{X}^{\operatorname{sm}}$ are finite and surjective.
If such $W$ exists, when $\dim(O_{\mathcal{E}}\cap M^{\operatorname{sm}}_{\mathscr{X}})\geq\mathrm{d}(\mathcal{E})-i=\frac{\mathbf{v}^{2}+2}{2}-i$, we can find a subvariety
$Z\subset O_{\mathcal{E}}\cap M^{\operatorname{sm}}_{\mathscr{X}}$
with $\dim Z\geq\mathrm{d}(\mathcal{E})-i$. Then $Z^{\prime}=\pi_{2}(\pi_{1}^{-1}(Z))$ is a constant cycle subvariety by Theorem 2.6 and we have $\dim Z^{\prime}=\dim Z\geq\frac{\mathbf{v}^{2}+2}{2}-i$. This means $\dim O_{E}\geq\frac{\mathbf{v}^{2}+2}{2}-i$ for any $E\in Z^{\prime}$. According to the proof of [15, Theorem 1.1], we have
$c_{2}(E)\in\mathbf{S}_{i}(X)$
where $E\in Z^{\prime}$. From the construction of $R$, we have $c_{2}(\mathcal{E})-c_{2}(E)\in\mathbf{S}_{0}(X)$ which implies $c_{2}(\mathcal{E})\in\mathbf{S}_{i}(X)$.
Proof of Claim 1
Let us explain the construction of $W$. Let $\Sigma_{\mathscr{X}}\subseteq M_{\mathscr{X}}$ and $\Sigma_{X}\subseteq M_{X}$ be a constant cycle Lagrangian subvariety defined in the proof of Proposition 3.2. Their restrictions to the smooth locus $M_{\mathscr{X}}^{\operatorname{sm}}$ (resp. $M_{X}^{\operatorname{sm}}$) are multisections of the abelian fibration of degree $d_{1}$ (resp. $d_{2}$).
Let $\Sigma_{\mathscr{X},C}$ and $\Sigma_{X,C}$ be their restriction to the fiber over $C\in|L|^{\operatorname{sm}}$. Denote by $i_{C}:C\hookrightarrow X$ the closed embedding. We can define
$W=\Big{\{}(\mathcal{E},E)\in M_{\mathscr{X}}^{\operatorname{sm}}\times M_{X}^{\operatorname{sm}}|~{}(i_{C})^{\ast}(\mathcal{E}-\frac{1}{d_{1}}\Sigma_{\mathscr{X},C})=(i_{C})^{\ast}(E-\frac{1}{d_{2}}\Sigma_{X,C})\in\operatorname{Pic}^{0}_{\mathbb{Q}}(C)\Big{\}}.$ (16)
By Example 2.4, for $(\mathcal{E},E)\in W$, we have
$c_{2}(\mathcal{E})-\frac{1}{d_{1}}c_{2}(\Sigma_{\mathscr{X},C})=c_{2}(E)-\frac{1}{d_{2}}c_{2}(\Sigma_{X,C})\in\operatorname{CH}_{0}(X).$ (17)
Here, $c_{2}(\Sigma_{\mathscr{X},C})=\sum\limits_{\mathcal{F}\in\Sigma_{\mathscr{X},C}}c_{2}(\mathcal{F})$ and $c_{2}(\Sigma_{X,C})=\sum\limits_{F\in\Sigma_{X,C}}c_{2}(F)$.
By Proposition 3.6, for every element $\mathcal{F}\in\Sigma_{\mathscr{X},C}$, we have $c_{2}(\mathcal{F})\in\mathbf{S}_{0}(X)$. Thus
$\frac{1}{d_{1}}c_{2}(\Sigma_{\mathscr{X},C})=\frac{1}{d_{2}}c_{2}(\Sigma_{\mathscr{X},C})\in\mathbf{S}_{0}(X).$
Here, the equality holds because they have the same degree. This forces $c_{2}(\mathcal{E})=c_{2}(E)$. Moreover, one can easily check that the projections are finite and surjective. Indeed, $W$ is obtained by taking the graph of the natural isomorphism between the two torsors $M_{\mathscr{X}}^{\operatorname{sm}}\to|L|^{\operatorname{sm}}$ and $M_{X}^{\operatorname{sm}}\to|L|^{\operatorname{sm}}$ after the finite base change
$\Sigma_{\mathscr{X}}^{{\operatorname{sm}}}\times_{|L|^{\operatorname{sm}}}\Sigma_{X}^{{\operatorname{sm}}}\to|L|^{{\operatorname{sm}}}.$
This completes the proof.
4. Twisted Beauville-Voisin class
In this section, we construct the twisted Beauville-Voisin class on $\mathscr{X}$.
4.1. Construction of $\mathfrak{o}_{\mathscr{X}}$
Throughout this section, all the Mukai vectors are assumed to be primitive and stability conditions are generic.
Definition 4.1.
Let $N=M_{\sigma}(\mathscr{X},\mathbf{w})$ be a Bridgeland moduli space and let $N^{\rm l.f.}\subseteq N$ be the open subset which consists of $\mathscr{X}$-twisted locally free sheaves. Let $\mathcal{F}\in N^{\rm l.f.}$ be a point lying on a constant cycle Lagrangian subvariety $Z$. We define
$\mathfrak{o}_{\mathscr{X}}=\frac{c_{2}(\mathcal{F})}{\operatorname{rank}\mathcal{F}}+(1-\frac{\deg(\mathcal{F})}{\operatorname{rank}\mathcal{F}})\mathfrak{o}_{X}.$ (18)
to be a zero cycle of degree $1$.
Remark 4.2.
The construction of $\mathfrak{o}_{\mathscr{X}}$ is clearly invariant under the action of tensoring line bundles. This is because $c_{2}(\mathcal{F})-c_{2}(\mathcal{F}\otimes L)\in\mathbf{S}_{0}(X)$.
The existence of such $N$ and $\mathcal{F}$ in ensured by the following result.
Proposition 4.3.
Let $\mathbf{v}\in\widetilde{\mathrm{H}}(\mathscr{X})$ and $\mathbf{v}^{\perp}\subseteq\widetilde{\mathrm{H}}_{\rm alg}(\mathscr{X})$ contains an isotropic element. Assume that $\operatorname{rank}(\mathbf{v})=r>\frac{\mathbf{v}^{2}+2}{2}$ and there exists a slope stable element in $M_{H}(\mathscr{X},\mathbf{v})$. Then $M_{H}(\mathscr{X},\mathbf{v})$ contains a constant cycle Lagrangian subvariety $Z$ such that $Z\cap M_{H}^{\rm l.f.}(\mathscr{X},\mathbf{v})\neq\emptyset$.
Proof.
By (9), we know that $\operatorname{NS}(M_{H}(\mathscr{X},\mathbf{v}))\cong\mathbf{v}^{\perp}$ contains an isotropic element. Then $M_{H}(\mathscr{X},\mathbf{v})$ admits a birational Lagrangian fibration by Theorem 2.7. By Proposition 3.2, there is a dense collection of constant cycle Lagrangian subvarieties on $M_{H}(\mathscr{X},\mathbf{v})$. So it suffices to show $M_{H}^{\rm l.f.}(\mathscr{X},\mathbf{v})\neq\emptyset$.
Note that under the condition $r>\frac{\mathbf{v}^{2}+2}{2}$, every $\mu_{H}$-stable $\mathscr{X}$-twisted sheaves are locally free (cf[27, Remark 3.2] and [11, Remark 6.1.9]) and the assertion follows from our assumption.
Remark 4.4.
The existence of slope stable elements can be easily achieved. Let $r_{0}$ be the minimal rank of $\mathscr{X}$-twisted sheaves in $\mathrm{Coh}^{(1)}(\mathscr{X})$. Then $r_{0}$ is the order of $[\mathscr{X}]$ in $\operatorname{Br}(X)$ and every element in $\mathrm{Coh}^{(1)}(\mathscr{X})$ has rank $ar_{0}$ for some $a\in\mathbb{Z}$ (cf. [28, Lemma 3.2]). For any primitive vector
$\mathbf{v}=(ar_{0},h,s)\in\widetilde{\mathrm{H}}(\mathscr{X})$
with $a>0$ and $\gcd(a,H\cdot h)=1$, we have $\mathcal{E}\in M_{H}(\mathscr{X},\mathbf{v})$ is $\mu_{H}$-stable. When $r>\frac{\mathbf{v}^{2}+2}{2}$, we actually have $M_{H}^{\rm l.f.}(\mathscr{X},\mathbf{v})=M_{H}(\mathscr{X},\mathbf{v})$.
Moreover, to see the existence of $\mathbf{v}$ whose orthogonal complement contains an isotropic element, one can just let $s=0$ and the isotropic vector $(1,0,0)$ is contained in $\mathbf{v}^{\perp}$.
Now we show that $\mathfrak{o}_{\mathscr{X}}$ is independent of the choice of the moduli space $N$ and the constant cycle Lagrangian subvariety $Z$. The key result is
Proposition 4.5.
Let $\mathscr{M}_{i}:=\mathscr{M}_{\sigma_{i}}(\mathscr{X},\mathbf{v}_{i}),~{}i=1,\ldots,k$ and $\mathscr{N}_{j}:=\mathscr{M}_{\tau_{j}}(\mathscr{X},\mathbf{w}_{j}),~{}j=1,\ldots,\ell$ be Bridgeland moduli stacks. Set $M_{i}$ (resp. $N_{j}$) to be the corresponding coarse moduli space. Assume that
  1. (i)
    $M_{i}^{\rm l.f.}$ and $N_{j}^{\rm l.f.}$ are non-empty;
  2. (ii)
    $\mathcal{E}_{i}^{\vee}\otimes\mathcal{F}_{j}$ is globally generated and
    $\mathrm{H}^{n}(X,\mathcal{E}_{i}^{\vee}\otimes\mathcal{F}_{j})=0,\quad\forall~{}n=1,2,$ (19)
    for any $\mathcal{E}_{i}\in M_{i}$ and $\mathcal{F}_{j}\in N_{j}$;
  3. (iii)
    for $\mathcal{G}\in M_{i}$ (or $N_{i}$) and $\mathcal{G}^{\prime}\in M_{i^{\prime}}$ (resp. $N_{i^{\prime}}$), we have
    $\mathrm{H}^{1}(\mathcal{G}^{\vee}\otimes\mathcal{G}^{\prime})=0,\quad\forall~{}i\neq i^{\prime}$ (20)
    and
    $\mathrm{H}^{0}(\mathcal{G}^{\vee}\otimes\mathcal{G}^{\prime})=0,\quad\forall~{}i>i^{\prime}$ (21)
Let $\mathcal{E}=\bigoplus\limits_{i=1}^{k}\mathcal{E}_{i}$ and $\mathcal{F}=\bigoplus\limits_{j=1}^{\ell}\mathcal{F}_{j}$ with $\mathcal{E}_{i}\in M_{i}^{\rm l.f.}$ and $\mathcal{F}_{j}\in N_{j}^{\rm l.f.}$. If $\operatorname{rank}\mathcal{E}=\operatorname{rank}\mathcal{F}$, then
$c_{2}(\mathcal{E})-c_{2}(\mathcal{F})\in\mathbf{S}_{d}(X).$ (22)
where $d=\sum\limits_{i=1}^{k}(\mathrm{d}(\mathcal{E}_{i})-n_{i})+\sum\limits_{j=1}^{\ell}(\mathrm{d}(\mathcal{F}_{j})-m_{j})$, $\dim O_{\mathcal{E}_{i}}\cap M_{i}^{\rm l.f.}=n_{i}$ and $\dim O_{\mathcal{F}_{j}}\cap N_{j}^{\rm l.f.}=m_{j}$.
The proof will be given later. With this result, we have
Corollary 4.6.
$\mathfrak{o}_{\mathscr{X}}$ is independent of the choice of $N$ and $\mathcal{F}$.
Proof.
First, we fix $\mathbf{v}=(r,L+rB,0)$ a primitive Mukai vector which satisfies the assumptions in Proposition 4.3. Let $\mathcal{F}$ be a point lying on a constant cycle Lagrangian subvariety $Z\subseteq N$.
Suppose $N^{\prime}$ is another moduli space of twisted locally free sheaves and it contains a constant cycle Lagrangian subvariety $Z^{\prime}$. Let $\mathcal{E}\in N^{\prime}$. We set $\ell=\operatorname{rank}(\mathcal{F})$ and $k=\operatorname{rank}(\mathcal{E})$. Then one can take an ample line bundle $H$ and define
  • $\mathcal{E}_{i}=\mathcal{E}\otimes H^{\otimes a_{i}},~{}i=1,\ldots,\ell$.
  • $\mathcal{F}_{j}=\mathcal{F}\otimes H^{\otimes b_{j}},~{}j=1,\ldots,k$;
By choosing $b_{j}-b_{j-1}$, $a_{i}-a_{i-1}$ and $b_{1}-a_{k}$ sufficiently large, $\mathcal{F}_{i}$ and $\mathcal{E}_{i}$ satisfy the conditions in Proposition 4.5 by Lemma 4.7 below. Due to Remark 4.2, this does not change the associated twisted BV class. Thus we have
$c_{2}(\bigoplus\limits_{j=1}^{k}\mathcal{F}_{i})-c_{2}(\bigoplus\limits_{i=1}^{\ell}\mathcal{E}_{i})\in\mathbf{S}_{0}(X)$ (23)
which is equivalent to $kc_{2}(\mathcal{F})-\ell c_{2}(\mathcal{E})\in\mathbf{S}_{0}(X)$. ∎
Lemma 4.7.
Let $M$ be a scheme of finite type. Let $\mathscr{E}$ be a coherent sheaf on $M\times X$ and let $H$ be an ample line bundle on $X$. Then there exists $d\in\mathbb{N}$ such that whenever $k\geq d$, $\mathrm{H}^{i}(\mathscr{E}_{b}\otimes H^{\otimes k})=0,i=1,2$ for any $b\in M$.
Proof.
This can be done by induction on the dimension of $M$. When $\dim M=0$, this is clear. Suppose this holds when $\dim M\leq n$. If $\dim M=n+1$, let $N^{i}$ be the support of the sheaf
$\mathbf{R}^{i}f_{*}(\mathscr{E}\otimes g^{*}H^{\otimes k}),~{}i=1,2$
which is a closed subscheme of $M$ of finite type. Here, $f:M\times X\to M$ and $g:M\times X\to X$ are projections. By Serre vanishing, there exists $d_{0}\in\mathbb{N}$ such that $\dim N^{i}<\dim M$ when $k>d_{0}$. By induction hypothesis, there exists $d_{1}\in\mathbb{N}$ such that $\mathrm{H}^{i}(\mathcal{E}_{b}\otimes H^{\otimes k})=0$ for $b\in N_{i}$. Take $d=\max\{d_{0},d_{1}\}$. ∎
Remark 4.8 (Descent property).
We speculate that $\mathfrak{o}_{\mathscr{X}}=\mathfrak{o}_{X}$. The main evidence is that due to our construction, the twisted BV class $\mathfrak{o}_{\mathscr{X}}$ has nice descent property. This follows from the descent property of constant cycle Lagrangian subvarieties (cf. [8, Proposition 3.6]).
Here we can give a more explicit proof. Suppose $X$ is defined over $\bar{k}$. Take a Mukai vector $\mathbf{v}=(r,h,s)$ such that $\mathbf{v}^{2}=0$ and $r>1$. Then $M=M_{H}(\mathscr{X},\mathbf{v})$ is a K3 surface defined over $\bar{k}$ and it consists of Gieseker stable $\mathscr{X}$-twisted locally free sheaves. Take a $\bar{k}$-rational curve $C\subseteq M$ and let $\mathcal{F}$ be a point on $C$. Then $\mathfrak{o}_{\mathscr{X}}=\frac{c_{2}(\mathcal{F})}{r}+(1-\frac{\deg(\mathcal{F})}{r})\mathfrak{o}_{X}$ is clearly defined over $\bar{k}$.
When $\bar{k}=\bar{\mathbb{Q}}$, the Bloch-Beilinson conjecture predicts $\operatorname{CH}_{0}(X_{\bar{\mathbb{Q}}})=\mathbb{Z}\mathfrak{o}_{X}$ and $\mathfrak{o}_{\mathscr{X}}=\mathfrak{o}_{X}$.
4.9. Proof of Proposition 4.5
The idea is to consider the degenerate loci of morphisms between twisted sheaves. Let us recall the definition and some basic properties.
Definition 4.2.
Let $\mathcal{E}$ and $\mathcal{F}$ be two $\mathscr{X}$-twisted locally free sheaves. For $t:\mathcal{E}\to\mathcal{F}\in\operatorname{Hom}(\mathcal{E},\mathcal{F})$, we define the $k$-th degenerate locus $\mathscr{D}_{k}(t)$ to be the closed substack of $\mathscr{X}$ defined by the image of the map
$\wedge^{k+1}t^{\vee}:\mathscr{H}om(\wedge^{k+1}\mathcal{E},\wedge^{k+1}\mathcal{F})^{\vee}\to\mathcal{O}_{\mathscr{X}}.$
We can also define the schematic $k$-th degenerate locus $D_{k}(t)$ which is the scheme-theoretic image in $X$. (cf[16, Definition 2.2.6.2])
Form the definition, since $\mathscr{H}om(\wedge^{k+1}\mathcal{E},\wedge^{k+1}\mathcal{F})$ is the pullback of a sheaf on $X$, $\mathscr{D}_{k}(t)\to D_{k}(t)$ is the pullback of $\mathscr{X}\to X$ via $D_{k}(t)\to X$. The following result is well-known for untwisted locally free sheaves (cf[6, §4.1][12, Lemma 2.14]).
Lemma 4.10.
Let $\mathcal{E}$ and $\mathcal{F}$ be two $\mathscr{X}$-twisted locally free sheaves of rank $e$ and $f$. Suppose $\mathcal{E}^{\vee}\otimes\mathcal{F}$ is globally generated. Then for general $t\in\operatorname{Hom}(\mathcal{E},\mathcal{F})=\mathrm{H}^{0}(\mathscr{X},\mathcal{E}^{\vee}\otimes\mathcal{F})$, $D_{k}(t)$ is either empty or has pure codimension $(e-k)(f-k)$. Moreover, the singular locus ${\rm Sing~{}}D_{k}(t)\subset D_{k-1}(t)$.
Proof.
When $\mathscr{X}\to X$ is trivial, this is exactly [12, Lemma 2.14]. In the case $\mathscr{X}\to X$ is essentially trivial, after taking a twisted line bundle $\mathcal{L}\in\operatorname{Pic}(\mathscr{X})$, we can identify $\operatorname{Hom}_{\mathcal{O}_{\mathscr{X}}}(\mathcal{E},\mathcal{F})$ with $\operatorname{Hom}_{\mathcal{O}_{X}}(\mathcal{E}\otimes\mathcal{L}^{\vee},\mathcal{F}\otimes\mathcal{L}^{\vee})$ and their schematic degenerate loci are the same. The assertion follows from the untwisted version.
In general, we let $Y\to X$ be a Severi-Brauer variety of $\mathscr{X}\to X$ with the Cartesian diagram
[…]
Then $\mathscr{Y}\to Y$ is essentially trivial. Set $\mathcal{E}^{\prime}=\varpi^{\ast}\mathcal{E}$ and $\mathcal{F}^{\prime}=\varpi^{\ast}\mathcal{F}$. Then we have
$\pi^{*}(\mathcal{E}^{\vee}\otimes\mathcal{F})\cong(\mathcal{E}^{\prime})^{\vee}\otimes\mathcal{F}^{\prime}$
as untwisted sheaves. By projection formula, there is an isomorphism
$\pi^{\ast}:\mathrm{H}^{0}(X,\mathcal{E}^{\vee}\otimes\mathcal{F})\to\mathrm{H}^{0}(Y,(\mathcal{E}^{\prime})^{\vee}\times\mathcal{F}^{\prime}).$ (24)
Since $\mathscr{Y}\to Y$ is essentially trivial, the assertion holds for $\pi^{\ast}t$ when $t$ is general. Note that $D_{k}(\pi^{\ast}t)=\pi^{-1}D_{k}(t)$ from the definition, we therefore get
$\operatorname{codim}D_{k}(t)=\operatorname{codim}D_{k}(\pi^{\ast}t)=(e-k)(f-k).$
and ${\rm Sing~{}}\pi^{-1}D_{k}(t)\subset\pi^{-1}(D_{k-1}(t))$. As $Y\to X$ is a $\mathbb{P}^{n}$-fibration, one must have ${\rm Sing}~{}\pi^{-1}D_{k}(t)=\pi^{-1}({\rm Sing}~{}D_{k}(t))$. It follows that $D_{k}(t)\subseteq D_{k-1}(t)$ as well. ∎
Let us come back to the proof of Proposition 4.5. From the vanishing condition (20) in assumption (iii), we have
$\dim_{\mathbb{C}}\operatorname{Ext}^{1}(\mathcal{E},\mathcal{E})=2\sum\limits_{i=1}^{k}\mathrm{d}(\mathcal{E}_{i}){\rm~{}and~{}}\dim_{\mathbb{C}}\operatorname{Ext}^{1}(\mathcal{F},\mathcal{F})=2\sum\limits_{j=1}^{\ell}\mathrm{d}(\mathcal{F}_{j}).$
By assumption (ii), for general $t\in\operatorname{Hom}(\mathcal{E},\mathcal{F}),$ we have a short exact sequence
[…]
According to Lemma 4.10, $\mathcal{Q}_{t}$ is supported on a smooth curve $\mathscr{C}_{t}$ and it is locally free of rank $1$ on its support.
Let $\mathscr{E}_{i}\in\mathrm{D}^{(-1,1)}(\mathscr{X}\times\mathscr{M}_{i})$ and $\mathscr{F}_{i}\in\mathrm{D}^{(-1,1)}(\mathscr{X}\times\mathscr{N}_{j})$ be the universal sheaf. Set
$\mathscr{G}:=\boxtimes_{i,j}(\mathscr{E}_{i}^{\vee}\boxtimes\mathscr{F}_{j})\in\mathrm{Coh}^{(-1,1,1)}(X\times\prod_{i=1}^{k}\mathscr{M}_{i}\times\prod_{j=1}^{\ell}\mathscr{N}_{j}).$
For simplicity, we assume that $\mathscr{M}_{i}=M_{i}$ and $\mathscr{N}_{j}=N_{j}$ are fine moduli spaces. Otherwise, one may either work with everything over stacks directly or pass to the associated Severi-Brauer varieties just as in the proof of [15, Theorem 3.1]. Take an irreducible component $Y_{i}\subseteq O_{\mathcal{E}_{i}}$ of dimension $n_{i}$ and $Z_{j}\subseteq O_{\mathcal{F}_{j}}$ of dimension $m_{j}$, we set
$Y=\prod_{i=1}^{k}Y_{i}~{}{\rm and}~{}Z=\prod_{j=1}^{\ell}Z_{j}.$
Then we can restrict $\pi_{\ast}(\mathscr{G})$ to $Y\times Z$ which is locally free by assumption (i). Define $\mathbf{P}=\mathbb{P}((\pi_{\ast}\mathscr{G})|_{Y\times Z})$ to be the associated projective bundle over $Y\times Z$. There is a diagram
[…]
Claim 2.
$\psi(\mathbf{P}_{y_{1},z_{1}})\cap\psi(\mathbf{P}_{y_{2},z_{2}})\neq\emptyset$ if and only if $(y_{1},z_{1})=(y_{2},z_{2})$. Moreover, the fiber of $\psi$ over $\mathcal{Q}_{t}$ has dimension
$\mathrm{d}(\mathcal{E})+\mathrm{d}(\mathcal{F})-\frac{\langle\mathbf{v}(\mathcal{E}),\mathbf{v}(\mathcal{E})\rangle}{2}-\frac{\langle\mathbf{v}(\mathcal{F}),\mathbf{v}(\mathcal{F})\rangle}{2}-2$ (26)
for general $t$.
Using (26), we can deduce that
$\dim\mathrm{Image}(\psi)$ $=\dim Y+\dim Z+h^{0}(\pi_{\ast}\mathcal{G}|_{Y\times Z})-1-\dim\hbox{fiber of $\psi$}$
$=\sum\limits_{i=1}^{k}n_{i}+\sum\limits_{j=1}^{\ell}m_{j}-\langle\mathbf{v}(\mathcal{E}),\mathbf{v}(\mathcal{F})\rangle-(\mathrm{d}(\mathcal{E})+\mathrm{d}(\mathcal{F})-\frac{\langle\mathbf{v}(\mathcal{E}),\mathbf{v}(\mathcal{E})\rangle}{2}-\frac{\langle\mathbf{v}(\mathcal{F}),\mathbf{v}(\mathcal{F})\rangle}{2}-1)$
$=\sum\limits_{i=1}^{k}n_{i}+\sum\limits_{j=1}^{\ell}m_{j}+\mathrm{d}(\mathcal{Q}_{t})-(\mathrm{d}(\mathcal{E})+\mathrm{d}(\mathcal{F}))$
$=\mathrm{d}(\mathcal{Q}_{t})-d.$
Due to [19, Theorem], $\mathrm{Image}(\psi)$ is a constant cycle subvariety. By Theorem 3.8, we get
$c_{2}(\mathcal{E})-c_{2}(\mathcal{F})=c_{2}(\mathcal{Q}_{t})\in\mathbf{S}_{d}(X).$ (27)
as desired.
It remains to prove Claim 2. Suppose there is another short exact sequence
$0\to\mathcal{E}^{\prime}\xrightarrow{t^{\prime}}\mathcal{F}^{\prime}\to\mathcal{Q}_{t^{\prime}}\to 0,$ (28)
with $\mathcal{F}^{\prime}=\bigoplus\limits_{j=1}^{\ell}\mathcal{F}_{j}^{\prime}$ and $\mathcal{Q}_{t^{\prime}}\cong\mathcal{Q}_{t}$. After applying the functor $\operatorname{Hom}(\mathcal{F}_{i},-)$ to (28) and (25), we get
[…]
Here the vanishing is guaranteed by (19) and $\mathcal{F}_{i}$ is a direct summand of $\mathcal{F}$. Therefore, we have
$\dim_{\mathbb{C}}\operatorname{Hom}(\mathcal{F}_{i},\mathcal{F}^{\prime})=\dim_{\mathbb{C}}\operatorname{Hom}(\mathcal{F}_{i},\mathcal{Q}_{t}^{\prime})=\dim_{\mathbb{C}}\operatorname{Hom}(\mathcal{F}_{i},\mathcal{Q}_{t^{\prime}}^{\prime})=\dim_{\mathbb{C}}\operatorname{Hom}(\mathcal{F}_{i},\mathcal{F}).$
Note that by (20) and (21), we have
$\dim_{\mathbb{C}}\operatorname{Hom}(\mathcal{F}_{i},\mathcal{F}_{j})=\dim_{\mathbb{C}}\operatorname{Hom}(\mathcal{F}_{i},\mathcal{F}_{j}^{\prime})=-\langle\mathbf{v}_{i},\mathbf{v}_{j}\rangle,\quad\forall j>i$
and
$\dim_{\mathbb{C}}\operatorname{Hom}(\mathcal{F}_{i},\mathcal{F}_{j})=\dim_{\mathbb{C}}\operatorname{Hom}(\mathcal{F}_{i},\mathcal{F}_{j}^{\prime})=0,\quad\forall j<i.$
Therefore, we obtain
$\dim_{\mathbb{C}}\operatorname{Hom}(\mathcal{F}_{i},\mathcal{F}_{i}^{\prime})=\dim_{\mathbb{C}}\operatorname{Hom}(\mathcal{F}_{i},\mathcal{F}_{i})=1.$
This forces $\mathcal{F}\cong\mathcal{F}^{\prime}$ since $\mathcal{F}_{i}$ and $\mathcal{F}_{i}^{\prime}$ are stable of the same phase. Similarly, if we apply the functor $\operatorname{Hom}(\mathcal{E}_{i},-)$, we can get
$\dim_{\mathbb{C}}\operatorname{Hom}(\mathcal{E}_{i}^{\prime},\mathcal{E}_{i})=\dim_{\mathbb{C}}\operatorname{Ext}^{2}(\mathcal{E}_{i},\mathcal{E}_{i}^{\prime})=\dim_{\mathbb{C}}\operatorname{Ext}^{2}(\mathcal{E}_{i},\mathcal{E}_{i})=1.$ (30)
This implies $\mathcal{E}\cong\mathcal{E}^{\prime}$ and proves the first claim. The fiber $\psi^{-1}(\mathcal{Q}_{t})$ consists of the kernel of $\mathcal{F}\to\mathcal{Q}_{t}$ and we have
$\dim\psi^{-1}(\mathcal{Q}_{t})$ $=\dim_{\mathbb{C}}\operatorname{Hom}(\mathcal{F},\mathcal{Q}_{t})+\dim_{\mathbb{C}}\operatorname{Hom}(\mathcal{E},\mathcal{E})-\dim_{\mathbb{C}}\operatorname{Hom}(\mathcal{Q}_{t},\mathcal{Q}_{t})-1$
$=\dim_{\mathbb{C}}\operatorname{Hom}(\mathcal{F},\mathcal{F})+\dim_{\mathbb{C}}\operatorname{Hom}(\mathcal{E},\mathcal{E})-\dim_{\mathbb{C}}\operatorname{Hom}(\mathcal{Q}_{t},\mathcal{Q}_{t})-1$
$=\mathrm{d}(\mathcal{F})-\frac{\langle\mathbf{v}(\mathcal{F}),\mathbf{v}(\mathcal{F})\rangle}{2}+\mathrm{d}(\mathcal{E})-\frac{\langle\mathbf{v}(\mathcal{E}),\mathbf{v}(\mathcal{E})\rangle}{2}-2.$
Here, $\dim_{\mathbb{C}}\operatorname{Hom}(\mathcal{F},\mathcal{Q}_{t})=\dim_{\mathbb{C}}\operatorname{Hom}(\mathcal{F},\mathcal{F})$ is obtained by applying the functor $\operatorname{Hom}(\mathcal{F},-)$ to (25) and $\dim_{\mathbb{C}}\operatorname{Hom}(\mathcal{Q}_{t},\mathcal{Q}_{t})=1$ as $\mathcal{Q}_{t}$ is stable. ∎
4.11. Proof of Theorem 1.2
Let $\mathcal{E}\in\mathrm{D}^{b}(\mathscr{X})$. As the argument proceeds a similar way as in [25, 21], we only sketch the proof.
i). $\mathcal{E}$ is a sheaf
Set $r=\operatorname{rank}(\mathcal{E})$. If $r=0$, this is Theorem 3.8. When $r>0$, we have the following situations
  • If $\mathcal{E}$ is locally free and simple, we can find two primitive Mukai vectors $\mathbf{v}^{\prime},\mathbf{w}\in\widetilde{\mathrm{H}}(\mathscr{X})$ satisfying the conditions in Proposition 4.3 and
    $\operatorname{rank}(\mathbf{v}^{\prime})+\operatorname{rank}(\mathcal{E})=\operatorname{rank}(\mathbf{w}).$
    The existence of such vectors is ensured by Remark 4.4. By Proposition 4.3, there exist $\mathscr{X}$-twisted locally free sheaves
    $\mathcal{E}^{\prime}\in M_{H}(\mathscr{X},\mathbf{v}^{\prime})$ and $\mathcal{F}\in M_{H}(\mathscr{X},\mathbf{w})$
    lying on some constant cycle Lagrangian subvarieties respectively. In other words, $\dim O_{\mathcal{E}^{\prime}}=\mathrm{d}(\mathcal{E}^{\prime})$ and $\dim O_{\mathcal{F}}=\mathrm{d}(\mathcal{F})$.
    As in the proof of Corollary 4.6, up to tensoring sufficiently ample line bundles, we assume $\mathcal{E}$, $M_{H}(\mathscr{X},\mathbf{v}^{\prime})$ and $M_{H}(\mathscr{X},\mathbf{w})$ satisfy the conditions in Proposition 4.5. It follows from Proposition 4.5 that
    $c_{2}(\mathcal{E})+c_{2}(\mathcal{E}^{\prime})-c_{2}(\mathcal{F})\in\mathbf{S}_{\mathrm{d}(\mathcal{E})}(X).$
    This implies $c_{2}(\mathcal{E})-r\mathfrak{o}_{\mathscr{X}}\in\mathbf{S}_{\mathrm{d}(\mathcal{E})}(X)$ from the definition of $\mathfrak{o}_{\mathscr{X}}$.
  • If $\mathcal{E}$ is only torsion-free, the double dual $\mathcal{E}^{\vee\vee}$ of $\mathcal{E}$ is locally free. There is a short exact sequence of sheaves
    $0\to\mathcal{E}\to\mathcal{E}^{\vee\vee}\to\mathcal{Q}\to 0$
    where $\mathcal{Q}$ is a 0-dimensional sheaf whose support is of length $\ell$. A direct calculation yields $\mathrm{d}(\mathcal{Q})>\ell$ and $c_{2}(\mathcal{Q})\in\mathbf{S}_{\ell}(X)\subseteq\mathbf{S}_{\mathrm{d}(\mathcal{Q})}(X)$. When $\mathcal{E}$ is slope stable, the double dual $\mathcal{E}^{\vee\vee}$ is also stable and hence simple. This gives
    $c_{2}(\mathcal{E}^{\vee\vee})-r\mathfrak{o}_{\mathscr{X}}\in\mathbf{S}_{\mathrm{d}(\mathcal{E}^{\vee\vee})}(X).$
    Consider the distinguished triangle $\mathcal{Q}[-1]\to\mathcal{E}\to\mathcal{E}^{\vee\vee}\to\mathcal{Q},$ we have
    $\operatorname{Hom}(\mathcal{Q}[-1],\mathcal{E}^{\vee\vee})=0,$
    as $\mathcal{Q}$ is 00-dimensional and $\mathcal{E}^{\vee\vee}$ is locally free. Then $\mathrm{d}(\mathcal{E})\leq\mathrm{d}(\mathcal{E}^{\vee\vee})+\mathrm{d}(\mathcal{Q}[-1])$. It follows that
    $c_{2}(\mathcal{E})-r\mathfrak{o}_{\mathscr{X}}=c_{2}(\mathcal{E}^{\vee\vee})-r\mathfrak{o}_{\mathscr{X}}-c_{2}(\mathcal{Q})\in\mathbf{S}_{\mathrm{d}(\mathcal{E}^{\vee})}(X)+\mathbf{S}_{\mathrm{d}(\mathcal{Q})}(X).$
    When $\mathcal{E}$ is not slope stable, one can obtain the assertion by applying the argument above to the HN factors of $\mathcal{E}$.
  • If $\mathcal{E}$ is not torsion free, there is a short exact sequence
    $0\to\mathcal{T}\to\mathcal{E}\to\mathcal{F}\to 0$ (31)
    with $\mathcal{T}$ torsion and $\mathcal{F}$ torsion-free. Then we have
    $c_{2}(\mathcal{E})-r\mathfrak{o}_{\mathscr{X}}$ $=c_{2}(\mathcal{T})+(c_{2}(\mathcal{F})-r\mathfrak{o}_{\mathscr{X}})$ (32)
    $\in\mathbf{S}_{\mathrm{d}(\mathcal{T})+\mathrm{d}(\mathcal{F})}(X)\subseteq\mathbf{S}_{\mathrm{d}(\mathcal{E})}(X).$
    The last inequality holds as $\operatorname{Hom}(\mathcal{T},\mathcal{F})=0$ and hence $\mathrm{d}(\mathcal{T})+\mathrm{d}(\mathcal{F})\leq\mathrm{d}(\mathcal{E})$.
ii). $\mathcal{E}$ is not a sheaf
There is a standard distinguished triangle
$\mathcal{F}\to\mathcal{E}\to\mathcal{G}\to\mathcal{F}[1].$
where $\mathcal{G}$ is the shifted sheaf $h^{m}(\mathcal{E})[-m]$ and $\mathcal{F}\in\mathrm{D}^{b}(X)$ is the truncated complex $\tau^{\leq m-1}$ By induction on the length of $\mathcal{E}$, one can get $c_{2}(\mathcal{E})-\operatorname{rank}(\mathcal{E})\mathfrak{o}_{\mathscr{X}}\in\mathbf{S}_{\mathrm{d}(\mathcal{E})}(X)$. ∎
4.12. SYZ filtration v.s. Voisin’s filtration
Now we can generalize the work in [21] and [15] to twisted K3 surfaces. With the twisted Beauville-Voisin class $\mathfrak{o}_{\mathscr{X}}$, we are allowed to define the twisted incidence correspondence
$\Gamma^{\prime}:=\Big{\{}(\xi,\mathcal{E})~{}|~{}c_{2}(\mathcal{E})-\operatorname{rank}(\mathcal{E})\mathfrak{o}_{\mathscr{X}}=[\xi]+k\mathfrak{o}_{X}\in\operatorname{CH}_{0}(X)\Big{\}}\subset X^{[d]}\times M_{\sigma}(\mathscr{X},\mathbf{v}),$ (33)
for arbitrary $\mathbf{v}$. As before, due to Theorem 1.2, the projection maps $\pi_{1}:\Gamma^{\prime}\to M_{\sigma}(\mathscr{X},\mathbf{v})$ and $\pi_{2}:\Gamma^{\prime}\to X^{[n]}$ are dominant.
So we can obtain a subvariety $\Gamma\subseteq\Gamma^{\prime}$ such that $\Gamma$ is generically finite over $M_{\sigma}(\mathscr{X},\mathbf{v})$ and $X^{[d]}$. For two general points $q,p$ in $\Gamma$ lying on the same fiber of $\pi_{1}$ (resp. $\pi_{2}$), we have $\pi_{2}(p)=\pi_{2}(q)\in\operatorname{CH}_{0}(X^{[d]})$ (resp. $\pi_{1}(p)=\pi_{1}(q)\in\operatorname{CH}_{0}(M_{\sigma}(\mathscr{X},\mathbf{v}))$) by Theorem 2.6. This implies that if $Z\subset X^{[d]}$ is a constant cycle subvariety, $\pi_{1}(\pi_{2}^{-1}(Z))$ is a constant cycle subvariety as well.
Theorem 4.13.
If $c_{2}(\mathcal{E})-\operatorname{rank}(\mathcal{E})\mathfrak{o}_{\mathscr{Y}}\in\mathbf{S}_{i}(X)$, then $\dim O_{\mathcal{E}}\geq d-i.$ In other words, there is an inclusion
$\mathbf{S}_{\bullet}^{\rm SYZ}\operatorname{CH}_{0}(M_{\sigma}(\mathscr{X},\mathbf{v}))\subseteq\mathbf{S}_{\bullet}^{\operatorname{BV}}\operatorname{CH}_{0}(M_{\sigma}(\mathscr{X},\mathbf{v})).$
Conversely, if $\dim O_{\mathcal{E}}\cap M^{\rm l.f.}_{\sigma}(\mathscr{X},\mathbf{v})\geq d-i$, then $\mathcal{E}\in\mathbf{S}_{i}(\mathrm{D}^{b}(\mathscr{X}))$. In particular, if $M_{\sigma}(\mathscr{X},\mathbf{v})$ consists of locally free sheaves, then
$\mathbf{S}_{\bullet}^{\rm SYZ}\operatorname{CH}_{0}(M_{\sigma}(\mathscr{X},\mathbf{v}))=\mathbf{S}_{\bullet}^{\operatorname{BV}}\operatorname{CH}_{0}(M_{\sigma}(\mathscr{X},\mathbf{v})).$
Proof.
The proof is identical to the proof of [21, Theorem 0.5]. ∎
As an application, we obtain a Zariski density result for constant cycle subvarieties.
Corollary 4.14.
For any $\mathcal{E}\in M^{\rm l.f.}_{\sigma}(\mathscr{X},\mathbf{v})$ with dim $\dim O_{\mathcal{E}}\cap M^{\rm l.f.}_{\sigma}(\mathscr{X},\mathbf{v})=d-i$, the subvarieties of the orbit $O_{\mathcal{E}}\cap M^{\rm l.f.}_{\sigma}(\mathscr{X},\mathbf{v})$ with dimension $d-i$ are dense in $M^{\rm l.f.}_{\sigma}(\mathscr{X},\mathbf{v})$.
Proof.
This can be obtained along the same lines in the proof [15, Corollary 3.2]. ∎
4.15. Further remarks
There is a natural (conjectural) O’Grady filtration for abelian surfaces, which induces a Shen-Yin-Zhao (SYZ) type filtration on albanese fiber of Bridgeland moduli space for (twisted) abelian surfaces. More precisely, if $A$ is projective abelian surface, an O’Grady type set-theoretic filtration is given by
$\mathbf{S}_{i}(A):=\mathbf{S}_{i}(\mathrm{Km}(A))\subseteq\operatorname{CH}_{0}(A)$ (34)
Here, $\operatorname{CH}_{0}(\mathrm{Km}(A))$ can be identified as a subset of $\operatorname{CH}_{0}(A)$ via the pullback of the rational map $A\dashrightarrow\mathrm{Km}(A)$. If $M=M_{\sigma}(\mathscr{A},\mathbf{v})$ is a smooth projective Bridgeland moduli space of $\sigma$-stable objects on a twisted abelian surface $\mathscr{A}\to A$, the kernel $K:=K_{\sigma}(\mathscr{A},\mathbf{v})$ of the albanese map
$\mathfrak{alb}\colon M_{\sigma}(\mathscr{A},\mathbf{v})$ $\longrightarrow\widehat{A}\times A,$ (35)
$\mathcal{F}$ $\mapsto(\det(\mathcal{F})\otimes\det(\mathcal{F}_{0})^{-1},\mathrm{alb}(c_{2}(\mathcal{F})-c_{2}(\mathcal{F}_{0})),$
is a smooth projective hyper-Kähler variety, where $\mathcal{F}_{0}\in M_{\sigma}(\mathscr{A},\mathbf{v})$. Note that for each $\mathcal{F}\in K$, one has
$c_{2}(\mathcal{F})-c_{2}(\mathcal{F}_{0})\in\operatorname{CH}_{0}(\mathrm{Km}(A))_{\hom}=\operatorname{CH}_{0}(A)_{\rm alb}.$
If there is a $\mathcal{F}_{0}$ with $c_{2}(\mathcal{F}_{0})\in\mathbf{S}_{0}(A)$, one can thus define a Shen-Yin-Zhao type filtration on $K_{\sigma}(\mathscr{A},\mathbf{v})$. In this situation, our construction of (twisted) incidence correspondences is also valid and one can obtain a similar result for the degenerate loci of stable locally free sheaves on twisted abelian surfaces. The main problem will be the very first step, that is to check whether the induced SYZ-type filtration coincide with Voisin’s filtration when $K=\ker(A^{[n+1]}\to A)$. Note that in this case, one does have an element $\mathcal{F}_{0}\in K$ with $c_{2}(\mathcal{F}_{0})\in\mathbf{S}_{0}(A)$.
5. Proof of Theorem 1.4
5.1. Beauville-Voisin subgroup
Let $\mathrm{R}(X)\subset\operatorname{CH}(X)$ be the Beauville-Voisin subring. Given a derived equivalence $\Psi:\mathrm{D}^{(1)}(\mathscr{X})\to\mathrm{D}^{(1)}(\mathscr{Y})$, it induces a (modified) additive group isomorphism
$\widetilde{\Psi}^{\operatorname{CH}}:\operatorname{CH}^{*}(X)$ $\to\operatorname{CH}^{*}(Y),$
$\widetilde{\operatorname{ch}}_{\mathscr{X}}(\mathcal{E})$ $\mapsto\widetilde{\operatorname{ch}}_{\mathscr{Y}}(\Psi(\mathcal{E}))Y$
where $\widetilde{\operatorname{ch}}_{\mathscr{X}}(\mathcal{E})=\operatorname{ch}_{\mathscr{X}}(\mathcal{E})-\operatorname{rank}(\mathcal{E})\mathfrak{o}_{\mathscr{X}}$ and $\widetilde{\operatorname{ch}}_{\mathscr{Y}}(\Psi(\mathcal{E}))=\operatorname{ch}_{\mathscr{Y}}(\Psi(\mathcal{E}))-\operatorname{rank}(\Psi(\mathcal{E}))\mathfrak{o}_{\mathscr{Y}}.$ Then the following result can be viewed as a generalization of [9].
Proposition 5.2.
The group isomorphism $\widetilde{\Psi}^{\operatorname{CH}}$ preserves the Beauville-Voisin group.
Proof.
Let $\mathbf{v}\in\widetilde{\mathrm{H}}(\mathscr{X})$ be a primitive Mukai vector whose orthogonal complement in $\widetilde{\mathrm{H}}_{\rm alg}(\mathscr{X})$ contains an isotropic element. Note that $\mathrm{R}(X)$ is spanned by such vectors. It suffices to show for some $\mathcal{E}\in M_{\sigma}(\mathscr{X},\mathbf{v})$ with $\widetilde{\operatorname{ch}}_{\mathscr{X}}(\mathcal{E})\in\mathrm{R}(X),$ we have
$\widetilde{\operatorname{ch}}_{\mathscr{Y}}(\Psi(\mathcal{E}))\in\mathrm{R}(Y).$ (36)
Let $\dim M_{\sigma}(\mathscr{X},v)=2n$. Note that $M_{\sigma}(\mathscr{X},\mathbf{v})$ admits a birational Lagrangian fibration. As proved in Proposition 3.2, there exists $\mathcal{E}\in M_{\sigma(\mathscr{X},\mathbf{v})}$ such that $O_{\mathcal{E}}$ has dimension $n$ and its maximal dimensional irreducible components are Zariski dense in $M_{\sigma}(\mathscr{X},\mathbf{v})$. Set $\mathcal{F}=\Psi(\mathcal{E})\in\mathrm{D}^{b}(\mathscr{Y})$ and $\mathbf{w}=\mathbf{v}(\mathcal{F})$. The derived equivalence $\Psi$ induces an isomorphism
$M_{\sigma}(\mathscr{X},\mathbf{v})$ $\to M_{\tau}(\mathscr{Y},\mathbf{w})$ (37)
$x$ $\mapsto\Psi(x)$
with $\tau=\Psi(\sigma)$. Then $\dim O_{\mathcal{F}}=n$ and its maximal dimensional irreducible components are also Zariski dense in $M_{\tau}(\mathscr{Y},\mathbf{w})$. Using the twisted incidence variety $\Gamma\subseteq M_{\tau}(\mathscr{Y},\mathbf{w})\times Y^{[n]}$ defined above, one can immediately get $c_{2}(\mathcal{F})-\operatorname{rank}(\mathcal{F})\mathfrak{o}_{\mathscr{Y}}\in\mathbf{S}_{0}(Y)$ as the projections $\Gamma\to Y^{[n]}$ and $\Gamma\to M_{\tau}(\mathscr{Y},\mathbf{w})$ are generically finite and dominant. ∎
Remark 5.3.
In [21, Question 4.1], it has been asked if the map
$\Psi^{\operatorname{CH}}:\operatorname{CH}^{*}(X)$ $\to\operatorname{CH}^{*}(Y),$
$\operatorname{ch}_{\mathscr{X}}(\mathcal{E})$ $\mapsto\operatorname{ch}_{\mathscr{Y}}(\Psi(\mathcal{E}))$
preserves the Beauville-Voisin group. By Proposition 5.2, this question is equivalent to ask whether $\mathfrak{o}_{\mathscr{X}}=\mathfrak{o}_{X}$.
5.4. Proof of Theorem 1.4
Suppose we have
$c_{2}(\mathcal{E})-\operatorname{rank}(\mathcal{E})\mathfrak{o}_{\mathscr{X}}=x_{1}+\ldots+x_{i}+k\mathfrak{o}_{\mathscr{X}},$
where $x_{1},\ldots,x_{i}$ are distinct points on $X$ and $k$ is a constant. Take the direct sum
$\mathcal{E}^{\prime}\cong\mathbb{C}_{x_{1}}\oplus\cdots\oplus\mathbb{C}_{x_{i}}\in\mathrm{D}^{b}(\mathscr{X}).$
Then $\widetilde{\operatorname{ch}}(\mathcal{E})-\widetilde{\operatorname{ch}}(\mathcal{E}^{\prime})\in\mathrm{R}(X)$. Since $\widetilde{\Psi}^{\operatorname{CH}}$ preserves the Beauville-Voisin group, we get
$\widetilde{\operatorname{ch}}(\Psi(\mathcal{E}))-\widetilde{\operatorname{ch}}(\Psi(\mathcal{E}^{\prime}))\in\mathrm{R}(Y).$ (38)
By Theorem 1.2, $\Psi(\mathcal{E}^{\prime})\in\mathbf{S}_{\mathrm{d}(\mathcal{E}^{\prime})}(\mathrm{D}^{b}(\mathscr{Y}))=\mathbf{S}_{i}(\mathrm{D}^{b}(\mathscr{Y}))$. Hence $\Psi(\mathcal{E})\in\mathbf{S}_{i}(\mathrm{D}^{b}(\mathscr{Y}))$ as well. ∎
6. Applications to Bloch’s conjecture
In this section, we consider the action of (anti)-symplectic birational automorphism on the Chow group of a Bridgeland moduli space.
6.1. Bloch’s conjecture for zero cycles
For a smooth projective hyper-Kähler variety $Y$, Beauville and Voisin predicted that there is an increasing filtration $\mathbf{S}_{\bullet}\operatorname{CH}^{\ast}(Y)$ which serves as an opposite filtration to the famous conjectural Bloch-Beilinson filtration. For zero cycles, it is expected
[…]
A more approachable question to examine these conjectures (and expectations) is to consider the action of (anti)-symplectic birational automorphisms on $\mathrm{Gr}_{\bullet}\operatorname{CH}_{0}(Y)$. More precisely, we have the following conjecture
Conjecture 6.2 (Bloch’s conjecture for Bridgeland moduli space).
Let $Y$ be a Bridgeland moduli space of dimension $2n$. Set
$\operatorname{Gr}_{s}\operatorname{CH}_{0}(Y):=\mathbf{S}_{s}^{\rm SYZ}\operatorname{CH}_{0}(Y)/\mathbf{S}_{s-1}^{\rm SYZ}\operatorname{CH}_{0}(Y)$
to be the $s$-th graded piece. For any $\phi\in\mathrm{Bir}(Y)$, we have
$\phi^{\ast}|_{\mathrm{H}^{2,0}(Y)}=\pm\operatorname{id}\Leftrightarrow\phi_{\ast}|_{\mathrm{Gr}_{s}\operatorname{CH}_{0}(Y)}=(\pm 1)^{s}\operatorname{id},~{}\forall~{}1\leq s\leq n.$
This conjecture has been studied in [14] for untwisted Bridgeland moduli spaces, i.e. when $Y$ is isomorphic to the moduli space of stable objects in $\mathrm{D}^{b}(X)$ for some K3 surface $X$. The method in [14] is also valid for twisted Bridgeland moduli spaces once there is a “right” formulation of $\mathbf{S}_{\bullet}^{\rm SYZ}\operatorname{CH}_{0}(Y)$. We will explain this in the next subsection.
6.3. Birational motives
We denote by $\mathfrak{h}^{\circ}(Y)$ the birational motive of $Y$ in the sense of [13]. It has been conjectured in [22] that there is a natural grading on $\mathfrak{h}^{\circ}(Y)$ compatible with the co-algebra structure on $\mathfrak{h}^{0}(Y)$ (cf[22, Conjecture 2]). When $Y=M_{\sigma}(X,\mathbf{v})$ for untwisted K3, such grading is governed by the isomorphism
$\mathfrak{h}^{\circ}(Y)\cong\mathfrak{h}^{\circ}(Y)_{(0)}\oplus\ldots\oplus\mathfrak{h}^{\circ}(Y)_{(n)}$
as co-algebra objects by [22, Theorem 4], where $\dim Y=2n$. The grading on $\mathfrak{h}^{\circ}(Y)$ is compatible with $\mathbf{S}_{\bullet}^{\rm SYZ}\operatorname{CH}_{0}(Y)$, i.e.
$\mathbf{S}_{i}^{\rm SYZ}\operatorname{CH}_{0}(Y)=\operatorname{CH}_{0}(\mathfrak{h}^{\circ}(Y)_{(0)}\oplus\ldots\oplus\mathfrak{h}^{\circ}(Y)_{(i)}),$
(cf[22, Section 6.4 and Theorem 7.3]). Due to Theorem 1.2, we can provide new examples for [22, Conjecture 2]. The idea is that the Shen–Yin–Zhao type filtration admits a “motivic” splitting on $\mathfrak{h}^{\circ}(Y)$.
Proposition 6.4.
If $Y=M_{\sigma}(\mathscr{X},\mathbf{v})$, then the twisted incidence variety $\Gamma$ induces an isomorphism
$\mathfrak{h}^{\circ}(M_{\sigma}(\mathscr{X},\mathbf{v}))\cong\mathfrak{h}^{\circ}(X^{[n]}),$
as co-algebra objects. Moreover, the isomorphism respects the natural gradings.
Proof.
As shown before, $\Gamma\subseteq Y\times X^{[n]}$ is a generically finite and dominant correspondence. The assertion is then a direct consequence of [22, Proposition 2.3]. ∎
The key result is
Lemma 6.5.
Let $Y=M_{\sigma}(\mathscr{X},\mathbf{v})$ as before. Let $Z\in\operatorname{CH}^{2n}(Y\times Y)_{\rm hom}$ be a correspondence. Then we have
$Z_{*}|_{\mathrm{Gr}_{s}\operatorname{CH}_{0}(Y)}=(\pm)^{s}\operatorname{id},$
if $Z_{*}|_{\mathrm{Gr}_{1}\operatorname{CH}_{0}(Y)}=\pm\operatorname{id}$.
Proof.
By Proposition 6.4, we only need to treat the case $M_{\sigma}(\mathscr{X},\mathbf{v})\cong X^{[n]}$. Then the result follows from [14, Proposition 3.8(iii)]. ∎
6.6. Proof of Theorem 1.8
By [1, Proposition 4][10, Lemma 2.6](see also [14, Corollary 3.3]), $Y\cong M_{\sigma}(\mathscr{X},\mathbf{v})$ for some twisted K3 surface $\mathscr{X}\to X$ and hence carries a Shen-Yin-Zhao type filtration on $\operatorname{CH}_{0}(Y)$.
One direction is well-known, see [24, Proposition 10.24] for symplectic case. Indeed, by [22, Lemma 2.1], if $\phi_{*}$ acts as ±$\pm\operatorname{id}$ on $\operatorname{Gr}_{1}\operatorname{CH}_{0}(Y)$, then it acts as ±$\pm\operatorname{id}$ on $\mathfrak{h}^{\circ}(Y)_{(1)}$. Since the cohomology of $\mathfrak{h}^{\circ}(Y)_{(1)}$ is the transcendental part $\mathrm{T}(X)\subset\mathrm{H}^{2}(X,\mathbb{Z})$, we get $\phi^{\ast}|_{\mathrm{T}(X)}=\pm\operatorname{id}$. It follows that $\phi^{\ast}|_{\mathrm{H}^{2,0}(X)}=\pm\operatorname{id}$.
For the converse, suppose $\phi\in\mathrm{Bir}(Y)$ is (anti-)symplectic. By [14, Theorem 2.17], we know that the action of $\phi$ on ${\mathrm{Gr}_{1}\operatorname{CH}_{0}(Y)}$ is ±$\pm\operatorname{id}$. The assertion then follows from Lemma 6.5. ∎